OpenStudy (dumbsearch2):

. Solve the following system using any method. 6x + 12y = 14 10x + 20y = 25 (6, 1) (2,2) (1, 1) No solution. The system is inconsistent. Infinite number of solutions. The system is dependent.

OpenStudy (dumbsearch2):

Thanks so much guysss !:)

OpenStudy (mathstudent55):

First, divide the first equation by 2 and the second equation by 5.

OpenStudy (dumbsearch2):

@HEEDCOM IN ADVANCE, silly! :-)

OpenStudy (mathstudent55):

6x + 12y = 14 10x + 20y = 25 3x + 6y = 7 2x + 4y = 5 Now multiply first equation by 2 and second equation by 3 and subtract the second equation from the first equation. 6x + 12y = 14 6x + 12y = 15 -------------(subtract) 0 + 0 = -1 0 = -1 is a false statement. That means the lines a parallel, and there are no solutions.

OpenStudy (dumbsearch2):

Thank you you are brillient!!! :) @mathstudent55